Ok, now I have no idea how to set this one up. I can solve it on my own I just need some help interpreting it. :(
The New River George Bridge, is 3030.5 feet long. It was completed in 1997 near Fayetteville, West Virginia, and is the world’s longest single arch steel span. Its total weight is 8.80 x 10^7 pounds and its heaviest single piece weight 1.84 x 10^5 pounds.
a) How many times greater is the total weight of the bridge than the weight of the heaviest single piece?

Answer by Theo(13342)

facts:
total weight is 8.80 * 10^7 pounds.
heaviest single piece weight is 1.84 * 10^5 pound.

total weight of the bridge is (8.80 * 10^7) / (1.84 * 10^5) times greater than the heaviest single piece.

(8.80 * 10^7) / (1.84 * 10^5) = (8.80/1.84) * (10^7/10^5) = (4.782608696 * 10^2) greater.

8.80 * 10^7 = 88,000,000

1.84 * 10^5 = 184,000

88,000,000 / 184,000 = 478.2608696 which is the same as:

4.782608696 * 10^2